OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4.

OpenStudy (anonymous):

@Michele_Laino @rational Would either of you be able to help me with this question

OpenStudy (michele_laino):

hint: given a subsequent parabola: \[x = a{y^2} + by + c\] the equation of the directrix, and the coordinate of its focus are: \[\begin{gathered} x = - \frac{{1 + {b^2} - 4ac}}{{4a}},\quad \left( {directrix} \right) \hfill \\ \hfill \\ F = \left( {\frac{{1 - {b^2} + 4ac}}{{4a}},\quad - \frac{b}{{2a}}} \right),\quad \left( {focus} \right) \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so we can write this algebraic system: \[\left\{ \begin{gathered} - \frac{{1 + {b^2} - 4ac}}{{4a}} = 4 \hfill \\ \hfill \\ \frac{{1 - {b^2} + 4ac}}{{4a}} = - 4 \hfill \\ \hfill \\ - \frac{b}{{2a}} = 0 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

from the third equation we get b=0

OpenStudy (michele_laino):

now, using b=0, we can rewrite the first two equations as below: \[\left\{ \begin{gathered} 1 - 4ac = - 16a \hfill \\ 1 + 4ac = - 16a \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

adding those two equation side by side, we get: 2=-32a so, what is a?

OpenStudy (michele_laino):

equations*